There was a pleasant surprise this morning when I arrived at the detector. We were in lock, and the nasty noise I mentioned yesterday had gone (although there was still quite a lot of noise making the data below ~100Hz rather jumpy). This noise source caused there to be huge wings/sidebands on the 120Hz and 180Hz lines in the data (these lines are harmonics of the 60Hz line, which comes from the AC frequency of the US mains electricity). The noise seemed to have started when one of the optical benches (on which sits a variety of optics, like mirrors, beam splitters and such) was "landed" - it had previously been floated on a very thin layer of air. So, one of the things that was done yesterday was to re-float the bench. It looks like it was this that improved things, but it's not entirely conclusive (other things like slight realignments were performed). None of that happened during my shift, but it was nice to see the work payoff.
During the shift we were still troubled by the low frequency anthropogenic noise and we only stayed in lock for about half the shift, but that's better than yesterday (when there had been issues other than seismic noise keeping us from reacquiring lock).
We had some excitement in the control room when we were invaded by a large party of local school kids, but the seemed reasonably restrained and asked some questions about the various monitors in the control room (the group was split into two to visit the control room and both groups asked about the two big clocks we have, one of which shows the local time [PDT at the moment] and the other that shows UTC). While they were there I had some pieces of paper, with complicated looking maths on it, so hopefully I looked like a proper scientist to the students.